# Towards Locally Deployable Fine-Tuned Causal Large Language Models for Mode Choice Behaviour

### v1

@@ 3.4 Reasoning Analysis: Topics, Factor Coverage, and Semantic Structure
For this analysis, we focus exclusively on the Swissmetro dataset. This decision is motivated by its extensive use in behavioural modelling and its wide adoption as a benchmark in recent studies exploring large language models (LLMs) for travel mode choice inference. Swissmetro provides a controlled setting with a clear set of alternatives and well-documented trade-offs, making it particularly suitable for examining how LLM-generated reasoning aligns with choice semantics. Restricting the analysis to this dataset allows for a deeper exploration of reasoning structures and their relationship with model performance, while minimising confounding factors introduced by differing dataset designs or context-specific variables.
The Swissmetro reasoning corpus resolves into a dominant family that weighs travel time against cost. After strict cleaning and BERTopic modelling, several numerical topics carry the same semantics; we therefore consolidate them under a single label (Time–Cost) and retain three smaller, semantically distinct themes: Income / Class, Convenience / Transfers, and a residual Artifact / Debug that captures prompt remnants and formatting noise. The consolidated geometry and the shot–type overlay appear in Figure 12. Points that lie close in the UMAP share vocabulary and phrase structure; separation indicates distinct justificatory semantics.
@@
Figure 12, left panel shows how a large, contiguous region corresponds to Time–Cost. Its footprint is not uniform: a dense central spine contains short, decisive sentences contrasting total times and costs; an upper arc contains longer phrases that balance time savings against cost differences and occasionally mention comfort or reliability; a lower arc aligns with formulations framed against the train alternative. These internal pockets explain why multiple BERTopic indices mapped to the same label: they represent sub-themes within one argumentative family rather than genuinely different topics. Two compact satellites sit adjacent to the main mass. Income / Class gathers references to income, first-class seating, and season-ticket affordability; lexical specificity pulls these statements together and away from the general trade-off cluster. Convenience / Transfers groups mentions of transfers, connections, and door-to-door practicality; the cluster remains small on Swissmetro and lies near the time–cost boundary because many texts still include an explicit time reference. The Artifact / Debug cluster sits well apart on the left, confirming that residual boilerplate has been successfully quarantined and does not contaminate the substantive clusters
@@
Factor coverage, measured by the five-factor Explanation Strength Index (ESI) over time, cost, comfort, convenience, frequency, shows systematic shifts with shot type (Figure 13). Zeroshot produces the highest means and the widest spreads: global mean ESI lies around 0.550.550.550.600.600.60, consistent with longer, list-like narratives that enumerate several factors. Few-shot prompts reduce ESI to approximately 0.400.400.400.430.430.43 on average and narrow the distributions, which reflects shorter, case-specific rationales once exemplars anchor the relevant trade-offs. The vertical reference lines in Figure 13 show that this shift holds across models, although the magnitude varies. Gemma 3 12B, Gemma 3 4B and Qwen 2.5 7B retain relatively high zeroshot ESI and then move to tighter, lower-ESI bands under few-shot while delivering strong accuracy. Stealth 1.2 7B exhibits comparatively low zeroshot ESI and the narrowest targeted curves; the model tends to issue compact, conservative justifications, which matches its calibrated probability profiles. Llama 3.2 1B shows the largest separation between zeroshot and few-shot curves and the lowest targeted mean, consistent with concise but often under-decisive statements.
@@ 3.5 Fine-tuned Model Evaluation and Comparison with Existing Literature and Methods
We fine-tuned Gemma 3 12B on Swissmetro and refer to the resulting model as “LiTransMC”. Its performance is summarized in Figure 15, which compares the model against the best untuned local baselines and published results on the same dataset, Swissmetro, from previous studies using proprietary large-scale models (Chat GPT) with different techniques to align LLM behaviour for mode choice experiments as well as discrete choice and machine learning modelling techniques for mode choice modelling on the same dataset.
On individual prediction, LiTransMC under targeted few‑shot attains a weighted F1 of 0.6845. This exceeds the best untuned local run (Gemma 3 12B, targeted few‑shot, 0.6246) and surpasses all GPT baselines from literature on the same dataset: GPT 3.5 Turbo at 0.648 (Mo et al., 2023b) and the GPT 4o variants in Liu et al. (2025b), namely zero‑shot 0.543 and few‑shot 0.594, as well as the same‑group persona variant at 0.657. It also outperforms classical models reported by Mo et al. (2023b) neural network 0.676, random forest 0.646, and MNL 0.639—and the MNL baseline in Liu et al. (2025b) at 0.606. Relative to the strongest prior large‑model result on this dataset, GPT 4o with persona inference and loading (0.683; (Liu et al., 2025b), LiTransMC is on par and even marginally higher while relying on a compact 12B parameter model and straightforward supervised adaptation rather than a two‑stage persona inference and loading pipeline. Beyond accuracy, locally hosted fine‑tuned models offer practical advantages: they run without external APIs, reduce operating cost, keep data on premises for privacy and safety, and make high‑quality behaviour modelling accessible to teams without access to proprietary frontier models.
On distributional alignment, LiTransMC achieves a JSD of 0.000245, which is lower than every comparator. It improves on the best untuned local JSD (0.000986 from a DeepSeek R1 Distill Llama run) by a factor of about four, and it reduces divergence by roughly eighty‑six times relative to Liu et al. (2025b) best GPT 4o configuration (0.021). The MNL baseline in same study records a JSD of 0.483, underscoring the magnitude of improvement in reproducing aggregate mode shares. Notably, even without fine‑tuning, the best local baselines already surpass GPT 4o in the zero‑shot and few‑shot settings on weighted F1, indicating that carefully prompted, locally hosted models can be competitive before any supervised adaptation.
@@ 4 Conclusions
This study systematically evaluated eleven locally deployable, open-access LLMs for transport mode choice prediction, establishing their viability and uncovering several critical insights. Model performance was found to hinge primarily on architecture and learning strategy, with variance decomposition showing these factors accounted for over 98% of the explained performance variance, while prompt engineering and temperature tuning offered only marginal, “last-mile” benefits. Among learning strategies, targeted few-shot prompting, where models are guided using contextually similar exemplars based on structured sample selection, emerged as particularly effective, consistently improving both accuracy and stability across datasets.
A central contribution of this research is the introduction of LiTransMC, the first fine-tuned causal LLM for mode choice prediction. Fine-tuned using a parameter efficient Q-LoRA approach, LiTransMC achieved state-of-the-art performance, attaining a weighted F1 score of 0.6845 and an exceptionally low Jensen-Shannon Divergence (0.000245), outperforming both untuned local models and even larger proprietary systems such as GPT 4o reported in prior studies as well as classical discrete choice and machine learning models. This fine-tuning process demonstrated that domain specific adaptation not only narrows the gap between individual prediction accuracy and aggregate distributional calibration but nearly eliminates it, enabling the generation of synthetic commuter agents that are individually accurate and collectively realistic, which is an essential requirement for robust transport policy modelling.
Our findings also delineate a clear hierarchy of model capabilities: (1) stable baselines (e.g., DeepSeek R1) that provide reliable, low-variance outputs ideal for initial deployment and prompt calibration; (2) sensitive high-performers (e.g., Gemma 3 12B) capable of peak accuracy but dependent on exemplar quality; and (3) rapid learners (e.g., Stealth 7B) that benefit most from carefully curated prompting. An effective deployment pathway thus involves the selection of the right model as per the availability of curated datasets and the complexity of the survey dataset.
Beyond predictive performance, the conversational paradigm adopted here provides an added dimension of cognitive transparency, the ability to generate and analyze natural-language reasoning aligned with behavioural theory. This dual capability of accurate prediction coupled with interpretable rationale distinguishes LLM based approaches from conventional ”black-box” AI models and equips policymakers with tools that are both analytically powerful and explainable.

### v2

@@ 3.3 Reasoning Analysis: Topics, Factor Coverage, and Semantic Structure
For this analysis, we focus exclusively on the Swissmetro dataset. This decision is motivated by its extensive use in behavioural modelling and its wide adoption as a benchmark in recent studies exploring LLMs for travel mode choice inference (Liu et al., 2025; Mo et al., 2023). Swissmetro provides a controlled setting with a clear set of alternatives and well-documented trade-offs, making it particularly suitable for examining how LLM-generated reasoning aligns with choice semantics.
The Swissmetro reasoning corpus resolves into a dominant family that weighs travel time against cost. After strict cleaning and BERTopic modelling, several numerical topics carry the same semantics; we therefore consolidate them under a single label (Time–Cost) and retain three smaller, semantically distinct themes: Income / Class, Convenience / Transfers, and a residual Artifact / Debug that captures prompt remnants and formatting noise. The consolidated geometry and the shot–type overlay appear in Figure 12. Points that lie close in the UMAP share vocabulary and phrase structure, while separation indicates distinct justificatory semantics.
@@
Figure 12, left panel shows how a large, contiguous region corresponds to Time–Cost. Its footprint is not uniform: a dense central spine contains short, decisive sentences contrasting total times and costs; an upper arc contains longer phrases that balance time savings against cost differences and occasionally mention comfort or reliability; a lower arc aligns with formulations framed against the train alternative. These internal pockets explain why multiple BERTopic indices mapped to the same label: they represent sub-themes within one argumentative family rather than genuinely different topics. Two compact satellites sit adjacent to the main mass. Income / Class gathers references to income, first-class seating, and season-ticket affordability; lexical specificity pulls these statements together and away from the general trade-off cluster. Convenience / Transfers groups mentions of transfers, connections, and door-to-door practicality; the cluster remains small on Swissmetro and lies near the time–cost boundary because many texts still include an explicit time reference. The Artifact / Debug cluster sits well apart on the left, confirming that residual boilerplate has been successfully quarantined and does not contaminate the substantive clusters.
@@
Factor coverage, measured by the five-factor Explanation Strength Index (ESI) over time, cost, comfort, convenience, frequency, shows systematic shifts with shot type (Figure 13). Zeroshot produces the highest means and the widest spreads: global mean ESI lies around 0.550.550.600.60, consistent with longer, list-like narratives that enumerate several factors. Few-shot prompts reduce ESI to approximately 0.400.400.430.43 on average and narrow the distributions, which reflects shorter, case-specific rationales once exemplars anchor the relevant trade-offs. The vertical reference lines in Figure 13 show that this shift holds across models, although the magnitude varies. Gemma 3 12B, Gemma 3 4B and Qwen 2.5 7B retain relatively high zeroshot ESI and then move to tighter, lower-ESI bands under few-shot while delivering strong accuracy. Stealth 1.2 7B exhibits comparatively low zeroshot ESI and the narrowest targeted curves; the model tends to issue compact, conservative justifications, which matches its calibrated probability profiles. Llama 3.2 1B shows the largest separation between zeroshot and few-shot curves and the lowest targeted mean, consistent with concise but often under-decisive statements.
@@ 3.4 Fine-tuned Model Evaluation and Comparison with Existing Literature and Methods
We fine-tuned Gemma 3 12B on Swissmetro and refer to the resulting model as “LiTransMC”. Its performance is summarized in Figure 15, which compares the model against the best untuned local baselines and published results on the same dataset, Swissmetro, from previous studies using proprietary large-scale models (Chat GPT) with different techniques to align LLM behaviour for mode choice experiments as well as discrete choice and machine learning modelling techniques for mode choice modelling on the same dataset using the same data split for training and testing as well as same evaluation metrics.
On individual prediction, LiTransMC under targeted few‑shot attains a weighted F1 of 0.6845. This exceeds the best untuned local run (Gemma 3 12B, targeted few‑shot, 0.6246) and surpasses all GPT baselines from literature on the same dataset: GPT 3.5 Turbo at 0.648 (Mo et al., 2023) and the GPT 4o variants in Liu et al. (2025), namely zero‑shot 0.543 and similarity based few‑shot 0.594, as well as the behavioural aligning efforts using targetted few shot with same‑group persona variant at 0.657. It also outperforms classical models reported by Mo et al. (2023) neural network 0.676, random forest 0.646, and MNL 0.639, and the MNL baseline in Liu et al. (2025) at 0.606. Relative to the strongest prior large‑model result on this dataset that uses GPT 4o with few shot examples utilizing persona inference and loading (0.683; (Liu et al., 2025)), LiTransMC is on par and even marginally higher while relying on a compact 12B parameter model and straightforward supervised adaptation rather than a two‑stage persona inference and loading pipeline. Beyond accuracy, locally hosted fine‑tuned models offer practical advantages: they run without external APIs, reduce operating cost, keep data on premises for privacy and safety, and make high‑quality behaviour modelling accessible to teams without access to proprietary frontier models.
On distributional alignment, LiTransMC achieves a JSD of 0.000245, which is lower than every comparator. It improves on the best untuned local JSD (0.000986 from a DeepSeek R1 Distill Llama run) by a factor of about four, and it reduces divergence by roughly eighty‑six times relative to Liu et al. (2025) best GPT 4o configuration (0.021). The MNL baseline in same study records a JSD of 0.483, underscoring the magnitude of improvement in reproducing aggregate mode shares. Notably, even without fine‑tuning, the best local baselines already surpass GPT 4o in the zero‑shot and few‑shot settings on weighted F1, indicating that carefully prompted, locally hosted models can be competitive before any supervised adaptation.
@@
4 Recommendations and Implications
Our experiments point to two primary factors for reliable mode–choice prediction with causal LLMs: (i) careful model selection and (ii) the choice of learning style. Variance decomposition shows that these two factors explain the vast majority of performance variation across datasets, while prompt format and decoding temperature contribute comparatively little. This implies that most of the attainable gains come from picking an appropriate model class and aligning the learning setup to the task, rather than from extensive prompt engineering or sampling stochasticity tuning.
On the learning style, moving from zero-shot to few-shot consistently improves accuracy, but uncurated examples increase run-to-run variability, which in turn will hinder the models reproducibility and reliability for forecasting tasks. Targeted, similarity-based few-shot learning both raises accuracy and narrows dispersion by retrieving demonstrations that are behaviourally close to the test instance using a multi-component similarity score over socio-demographics, trip attributes, and additional context, rather than relying on cosine similarity alone in matching the most relevant examples for few-shot prompting. In practice, this stabilizes predictions while improving mean performance and should be the default few-shot strategy for travel behaviour modelling.
Prompt engineering matters, but more as finishing steps than primary drivers. Across models and datasets, direct prompts generally outperform CoT/ReAct once demonstrations are available, and a conservative temperature performs best; raising temperature tends to yield negligible or negative effects. A simple two-step procedure works well in practice: first, fix a low temperature and choose between direct versus reasoning-style prompts based on validation; second, only then probe minor temperature changes if needed.
For the researcher community, these findings suggest the following practical workflow. Start by screening a small set of compact, open models that are known to be strong for this task family, and evaluate them under your target learning style; This step will dominate your returns. Then enable targeted few-shot using structured similarity over the variables that drive behavioural substitution in your context (for example, travel time and cost in the numeric component, exact matching for nominal attributes, and ordinal proximity for ordered factors). This similarity formulation proved more behaviourally aligned than embedding-based cosine approaches on structured survey data. Researchers could also experiment with the different weights attributed to the different components of the similarity score based on their domain knowledge and population heterogeneity. It is also recommended to use direct prompting at low temperature as your initial default, adjusting only at the last stage if you see robust validation gains.
For the practitioners, two deployment implications stand out. First, revealed-preference settings provide stronger baselines and more stable per-trip predictions than stated-preference settings; when possible, prime the model with RP exemplars and use SP data as an auxiliary source with targeted few-shot to counter hypothetical bias. Second, compact open models that are lightly fine-tuned on domain data can match or exceed proprietary APIs while preserving privacy and controlling cost through local inference, making routine behavioural analytics more accessible.
When fine-tuning is feasible, a parameter-efficient pass on an appropriate open model shifts the frontier outward: instant-level accuracy improves beyond strong untuned baselines and aggregate alignment tightens substantially, while eliminating dependence on external APIs. This makes a fine-tuned, locally hosted classifier a practical default for agencies and consultancies that must handle sensitive travel data.
The reasoning analysis adds important operational lessons. Targeted few-shot tightened explanations around the dominant time–cost trade-off while preserving a purposeful role for secondary cues such as transfers or ticketing, which helped protect minority mode probability mass. We therefore recommend reporting a compact reasoning card alongside predictions during development, then constraining explanations in production to a brief statement of the top drivers while archiving full rationale, prompt and exemplars for audit. Future work can formalize simple guardrails on explanation quality by setting thresholds on our reasoning diagnostics, for example, an Explanation Strength Index above 0.60, with runs flagged when this is violated. In parallel, researchers can test sensitivity to exemplar content by adding or removing specific cues such as transfers or ticketing and then measuring the resulting shifts in explanation topics and aggregate mode shares.
More broadly, this work aligns with the evolution towards the broader framework of agentic, human-in-the-loop transportation analytics. A specialized mode-choice LLM can act as a “choice agent” within an agentic transportation system: retrieving behaviourally similar exemplars, producing decisions with calibrated probabilities, explaining trade-offs in plain language, and handing off to downstream simulation or optimization tools for verification. The LiTransMC pipeline demonstrates that such specialized, privacy-preserving models are viable today and can serve as foundational building blocks for more interactive, human-centred planning workflows.
@@ 5 Conclusions
This study systematically evaluated eleven locally deployable, open-access LLMs for transport mode choice prediction and showed that reliable performance depends mainly on two design decisions: the choice of base model and the learning style. Prompt template and decoding temperature offered only marginal, last-mile effects relative to those primary levers. Moving from zero-shot to curated few-shot improved accuracy and stability across datasets, and a similarity-based approach to exemplar selection was consistently beneficial.
A central contribution is LiTransMC, a domain-adapted causal LLM fine-tuned with a parameter-efficient strategy. LiTransMC achieved a weighted F1 of 0.6845 and a Jensen–Shannon divergence of 0.000245, outperforming untuned local baselines and larger proprietary references reported in prior work, while narrowing the gap between instance-level accuracy and aggregate calibration. The results indicate that specialized, locally hosted models can meet or exceed the performance of general-purpose systems while preserving privacy and controlling cost.
Across datasets, a clear pattern emerged: models that are strong in zero-shot provide stable baselines, models that benefit most from high-quality demonstrations reach the highest peaks under curated few-shot, and compact learners improve markedly when examples are well matched. Revealed-preference data produced stronger baselines than stated-preference data, which suggests that empirically grounded decisions provide a cleaner learning signal; this points to opportunities to reduce hypothetical bias in future survey designs.
@@ Acknowledgements
Declaration of generative AI use
Locally deployed large language models were the subject of investigation in this paper using a methodology, protocol, and analysis entirely developed by the authors. AI use was confined to the research experiment as described in the manuscript. For manuscript preparation, AI tools (e.g., Grammarly, ChatGPT) were used only for grammar and language refinement, with no AI-generated original content. The authors reviewed and edited the content as needed and take full responsibility for the content of the published article.